OpenStudy (anonymous):

Complete the equation for the linear function whose graph contains the points (-1,-2) and (3,10). y+blank=blank(x+1)

OpenStudy (anonymous):

@paki @iGreen @itsbribro

OpenStudy (anonymous):

Hiya :)

OpenStudy (itsbribro):

heyy (:

OpenStudy (anonymous):

Ok so usually I don't do this but, I just don't get this question. My bloody teachers are useless.

OpenStudy (itsbribro):

lol its okay but i just logged in awhile ago and im still kinda sleepy lol so maybe i can tag someone

OpenStudy (anonymous):

Yea thats Cool :)

OpenStudy (itsbribro):

@LunyMoony @HaleyJlynn @helpmath123

OpenStudy (anonymous):

Hiyas :)

OpenStudy (anonymous):

um y aim kinda terrible at this kind of math lol

OpenStudy (anonymous):

LOL Thats kool :)

OpenStudy (lunymoony):

writing my graph

OpenStudy (anonymous):

Ok I just made my grpah

OpenStudy (lunymoony):

probably better than mine, lets see it

OpenStudy (anonymous):

Ok so technically I didn't make one . I printed one .

OpenStudy (lunymoony):

(-1,-2) and (3,10). y+blank=blank(x+1) = -1y -2 = 3x +10 this seem right @itsbribro

OpenStudy (lunymoony):

If so then its .... -1y = 3x +8 -3x - 1y = +8 -3x = 8 + (1)y 3x = -8 -y x= -8 -y/3

OpenStudy (anonymous):

Thanks Sooooooooo Much :) I am virtual giving you thousands of hugs :)

OpenStudy (lunymoony):

lol, np!

OpenStudy (anonymous):

If you ever need help In ELA or SS just ask me . I am fabulous at both.